OpenStudy (anonymous):

In the neighborhood park there is a circular grassy picnic area, but half of the grass has died. The city wants to replant the grass, but needs to know the square footage of the damaged area. If the circular picnic area is 150 feet across, what is the square footage of the damaged area? Let π = 3.14

OpenStudy (anonymous):

If the diameter is 150, then divide it to get the radius, which is 75. Then plug it into the equation A=(pi)(r^2) (75)^2=5625 5625*3.14=17662.5 That's the area of the entire grassy circle. We only need to find half of it. 8831.24ft. That's your final answer.